In the simple Keynesian model, an increase of one dollar in autonomous expenditure will cause equilibrium income to increase by a multiple of this one dollar increase. Explain the process by which this happens using different approached of explanations


1
Expert's answer
2021-01-07T08:29:38-0500

Answer:  

Simple keynesian model

Firm are assumed to make no tax payments, all taxes are paid by households. The central proposition of the simple keynesian model is that nation output (income) reaches its equilibrium value when output is equal to aggregate demands..... AD=AS=Y

AD=C+I, AS=C+S

1) circular flow


Circular flow diagram of income and output for a three sector economy.

Y=E-(1)

Y=GDP, E= Aggregate demand

2) "E" has three component------a)household

b) Business investment c) government demand for goods and services

Y=E=C+I+G, GDP=aggregate = government demand

A)


B)


An increase of on dollar in autonomous expenditure will cause equilibrium income to increase by a multiple of this one dollar increase...... S=I

K=multiplier, "\Delta Y"change in income,

"\Delta I" change in investment

"=\Delta Y=\Delta C+\Delta I, K=\Delta Y/\Delta I"

The working of the multiplier assumes the following......

Change in investment

|

Change income

|

Change in consumption

|

Change in income

"\Delta \Iota \to \Delta Y\to\Delta C\to\Delta Y"

Positive relation between investment and income

So it is prove that increase one dollar...... Increase income